OpenStudy (anonymous):

okay my daughter is trying to figure out her math she is a freshman and don't understand it if someone can please explain to her how to do it it is match each equation to its slope and y -intercept

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

do u have an example of an equation ?

OpenStudy (anonymous):

yes y = -3/4x +5

OpenStudy (anonymous):

The slope is -3/4 and the y-intercept is 5

OpenStudy (anonymous):

or x-y=6

OpenStudy (anonymous):

ok. so the y-intercept is 5. the slope is -3/4.

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=Zj-9Zc9yKYU

OpenStudy (anonymous):

x-y=6 can be rearranged as y=x-6, so the slope is 1 and the y-intercept is -6

OpenStudy (anonymous):

so it basically means that u start on positive 5 on the y-axis to graph. From there, you would go down 3 over 4. rise over run. it has a negative slope.

OpenStudy (anonymous):

does it help, martyden?

OpenStudy (anonymous):

i still dont understand .

OpenStudy (anonymous):

ok. would it help for me to draw it?

OpenStudy (saifoo.khan):

ok, my turn now. Where are you stuck?

OpenStudy (anonymous):

-4x-y=1

OpenStudy (anonymous):

ok. so first step is move the -4x over. What u really want to do is isolate the y. so -y=1+4x, divide by -1 to get y alone. so y=-4x-1.

OpenStudy (anonymous):

the slope is -4. the y-int is -1.

OpenStudy (saifoo.khan):

Loook now.

OpenStudy (anonymous):

would it help to graph it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

20x=4y-4

OpenStudy (saifoo.khan):

To graph, remember: when x intercpts y = 0 when y intercpts x = 0 Substitute to solve: the equation you have is: -4x-y=1

OpenStudy (anonymous):

-4y=-20x-4. get the y alone by dividing by -4. so y=5x+1

OpenStudy (anonymous):

the y-int is 1. the slope is 5.

OpenStudy (anonymous):

-7x+y=8

OpenStudy (anonymous):

get the y alone. just add 7x. y=7x+8 slope is 7 y int is 8

OpenStudy (saifoo.khan):

when y-intercpts 0 = 0 -4x-y=1 -4(0)-y = 1 -y = 1 y = 1 Now the Line will cut on y-axis @ 1. same as with x. when x interpts y = 0 -4x-y=1 -4x -(0) = 1 -4x = 1 x = -1/4 The line will cut the x axis on -1/4

OpenStudy (anonymous):

-x+y=13

OpenStudy (anonymous):

to isolate y, add x. y=x+13. the y int is 13 slope is 1

OpenStudy (saifoo.khan):



OpenStudy (anonymous):

7x=4y+25

OpenStudy (saifoo.khan):

PLEASE ma'am, Post one your problems one by one to avoid confusion for youself.

OpenStudy (saifoo.khan):

yourself*

OpenStudy (anonymous):

Im Not Confused .

OpenStudy (saifoo.khan):

LOL, then it's ok.

OpenStudy (saifoo.khan):

To graph, remember: when x intercpts y = 0 when y intercpts x = 0

OpenStudy (anonymous):

to isolate y, u need to move 4y and 7x. so -4y=-7x+25. y=7/4x-25/4

OpenStudy (saifoo.khan):

you just want to grapph this out?

OpenStudy (anonymous):

i already asked her that

OpenStudy (anonymous):

No I Dont Need A Graph .

OpenStudy (anonymous):

So What Does The Slope Equal ?

OpenStudy (saifoo.khan):

Slope = 7/4

OpenStudy (anonymous):

well basically, what you really need to remember is to isolate y to obtain an equation y=mx+b. the # in front of x is your slope.

OpenStudy (anonymous):

the b value is your y-int

OpenStudy (anonymous):

Can You Reduce 7 / 4 ?

OpenStudy (saifoo.khan):

Never.